# The theory of implicit operations

### v1

@@ 12. Structure theory
By symmetry it suffices to show that the statement holds for the case where $\phi=\mathsf{Ker}(g)$. Since $g\colon\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}\to\bm{C}$ is a homomorphism with $\bm{C}\in\mathsf{K}$, we have $\mathsf{Ker}(g)\in\mathsf{Con}_{\mathsf{K}}(\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*})$. Moreover, recall that $\bm{C}\in\textsf{K}_{\textsc{rfsi}}$ and that $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras by assumption. Therefore, from $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}/\mathsf{Ker}(g)\in\mathbb{I}\mathbb{S}(\bm{C})$ it follows that $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}/\mathsf{Ker}(g)$ is either trivial or belongs to $\mathsf{K}_{\textsc{rfsi}}$. Lastly, recall that $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}\leqslant\bm{A}\times\bm{A}$ is a subdirect product with $\bm{A}\in\mathsf{K}$. As $\mathsf{K}$ is a relatively congruence distributive quasivariety by assumption, we can apply 2.18, obtaining the desired conclusion. ∎
By symmetry and 12.16 we may assume that there exist $\eta_{1},\eta_{2}\in\mathsf{Con}_{\mathsf{K}}(\bm{A})$ such that $\mathsf{Ker}(g)=(\eta_{1}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$ and $\mathsf{Ker}(h)\in\{(\eta_{2}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)},(A^{2}\times\eta_{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}\}$. We have two cases depending on whether $\mathsf{Ker}(h)$ is $(\eta_{2}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$ or $(A^{2}\times\eta_{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$.
@@
Now, recall that $\langle a,b\rangle\in\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)$. By 12.17 we obtain $\langle a,b\rangle\in\eta_{1}$. Since $b\in A$, this implies $\langle\langle a,b\rangle,\langle b,b\rangle\rangle\in(\eta_{1}\times A^{2})$. On the other hand, from $a,b\in A$ and $\eta_{2}\in\mathsf{Con}(\bm{A})$ it follows that $\langle\langle a,b\rangle,\langle b,b\rangle\rangle\in(A^{2}\times\eta_{2})$. As $\langle a,b\rangle,\langle b,b\rangle\in\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)$, we obtain
@@
We begin by showing that $\mathsf{K}$ is a variety. As $\mathsf{K}$ is a quasivariety, by 2.5 it suffices to show that $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A})$ for every $\bm{A}\in\mathsf{K}$. To this end, consider $\bm{A}\in\mathsf{K}$ and $\theta\in\mathsf{Con}(\bm{A})$. We will show that $\theta=\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(\theta)$ which, in turns, implies $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{A})$, as desired. In view of 2.8, the congruence $\theta$ is the universe of a subalgebra $\bm{B}\leqslant\bm{A}\times\bm{A}$. Furthermore, for every $\langle a,b\rangle\in B=\theta$ we have $\langle a,a\rangle,\langle b,b\rangle\in\theta=B$ because $\theta$ is a congruence of $\bm{A}$. Therefore, we can apply 12.15, obtaining
@@
Let $\mathsf{M}$ be a congruence preserving Beth companion of $\mathsf{K}$. Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ and $\mathsf{K}$ is a quasivariety by assumption, we obtain that $\mathsf{M}$ is also a quasivariety (see 10.3(ii)). In addition, as $\mathsf{M}$ is a Beth companion of $\mathsf{K}$ and $\mathsf{K}$ is a quasivariety, $\mathsf{M}$ has the strong epimorphism surjectivity property by 11.6. From the assumption that $\mathsf{K}$ is relatively congruence distributive and 12.8 it follows that $\mathsf{M}$ is relatively congruence distributive as well. Lastly, we will prove that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras. Consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}_{\textsc{rfsi}}$ with $\bm{A}$ nontrivial. By 12.9 we have $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$. Furthermore, $\bm{A}\leqslant\bm{B}$ implies $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{A}$ is nontrivial, so is $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$ and the assumption that $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras guarantee that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$. Together with $\bm{A}\in\mathsf{M}$, this allows us to apply 12.9, obtaining $\bm{A}\in\mathsf{M}_{\textsc{rfsi}}$. Hence, we conclude that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras.
Therefore, $\mathsf{M}$ is a quasivariety with the strong epimorphism surjectivity property that, moreover, is relatively congruence distributive and such that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras. Thus, from 12.19 it follows that $\mathsf{M}$ is an arithmetical variety with the congruence extension property. Since $\mathsf{M}$ is a variety, we obtain $\mathsf{M}_{\textsc{fsi}}=\mathsf{M}_{\textsc{rfsi}}$. As $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras, we conclude that so is $\mathsf{M}_{\textsc{fsi}}$. ∎
@@
Let $\mathsf{K}$ be a relatively filtral quasivariety with the strong epimorphism surjectivity property. As $\mathsf{K}$ is relatively filtral, it is relatively congruence distributive and $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras (see, e.g., [24, Cor. 6.5(i, iv)]). Therefore, from 12.19 it follows that $\mathsf{K}$ is an arithmetical variety. Since $\mathsf{K}$ is a relatively filtral quasivariety, this yields that $\mathsf{K}$ is a congruence permutable filtral variety. As congruence permutable filtral varieties coincide with discriminator varieties (see, e.g., [18, 52]), we conclude that $\mathsf{K}$ is a discriminator variety. ∎
Furthermore, we recall that a join semilattice $\langle A;\lor\rangle$ is said to be dually Brouwerian when for all $a,b\in A$ there exists the smallest $c\in A$ such that $a\leqslant b\lor c$. Moreover, an element $a$ of a complete lattice $\bm{A}$ is compact when for every $X\subseteq A$ such that $a\leqslant\bigvee X$ there exists a finite $Y\subseteq X$ such that $a\leqslant\bigvee Y$. With every quasivariety $\mathsf{K}$ and algebra $\bm{A}$ we associate a join semilattice $\mathsf{Comp}_{\mathsf{K}}(\bm{A})$ whose universe is the set of compact elements of $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ and whose join operation ++ is defined by the rule $\theta+\phi=\mathsf{Cg}^{\bm{A}}_{\mathsf{K}}(\theta\cup\phi)$. Lastly, a member $\bm{A}$ of a quasivariety $\mathsf{K}$ is said to be simple relative to $\mathsf{K}$ when $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ has exactly two elements, and we say that $\mathsf{K}$ is relatively semisimple when every member of $\mathsf{K}_{\textsc{rsi}}$ is simple relative to $\mathsf{K}$. When $\mathsf{K}$ is a variety, we simply say that $\bm{A}$ is simple and $\mathsf{K}$ semisimple. We will rely on the fact that a quasivariety $\mathsf{K}$ is relatively filtral if and only if it is relatively semisimple and $\mathsf{Comp}_{\mathsf{K}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{K}$ (see [24, Thm. 6.3] and [78, Thms. 5 & 8]). [^9] This description of filtrality originated in the context of varieties (see [51, 52]) and was later extended to quasivarieties in [24]. Bearing this in mind, we will now prove 12.11.
@@
Let $\mathsf{M}$ be a Beth companion of a relatively filtral quasivariety $\mathsf{K}$. Since every relatively filtral quasivariety has the relative congruence extension property (see, e.g., [24, Cor. 6.5(i)]), so does $\mathsf{K}$. As $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, we conclude that $\mathsf{M}$ is congruence preserving by 12.4(ii).
We will show that $\mathsf{M}$ is also a relatively filtral quasivariety. The fact that $\mathsf{M}$ is a quasivariety is a consequence of Theorem 10.3(ii) and the assumption that $\mathsf{M}$ is a pp expansion of the quasivariety $\mathsf{K}$. Therefore, it only remains to prove that $\mathsf{M}$ is relatively filtral, i.e., that it is relatively semisimple and $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{M}$. To show that $\mathsf{M}$ is relatively semisimple, consider $\bm{A}\in\mathsf{M}_{\textsc{rsi}}$. Since $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$, we can apply 12.9, obtaining $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rsi}}$. As the quasivariety $\mathsf{K}$ is relatively semisimple (because it is relatively filtral by assumption), we obtain that $\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ has exactly two elements. Moreover, $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ because $\bm{A}\in\mathsf{M}$ and $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$. Therefore, $\mathsf{Con}_{\mathsf{M}}(\bm{A})$ has exactly two elements, whence $\bm{A}$ is simple relative to $\mathsf{M}$. Hence, we conclude that $\mathsf{M}$ is relatively semisimple, as desired. Next we prove that $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{M}$. Consider $\bm{A}\in\mathsf{M}$ and observe that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ by 10.2. Since $\mathsf{M}$ is congruence preserving, we have $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$, whence $\mathsf{Comp}_{\mathsf{M}}(\bm{A})=\mathsf{Comp}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\mathsf{Comp}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ is dually Brouwerian (because $\mathsf{K}$ is relatively filtral and $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$), we conclude that so is $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$. Thus, we conclude that $\mathsf{M}$ is a relatively filtral quasivariety.
Lastly, observe that $\mathsf{M}$ has the strong epimorphism surjectivity property because it is a Beth companion of the quasivariety $\mathsf{K}$ (see Theorem 11.6). Thus, $\mathsf{M}$ is a relatively filtral quasivariety with the strong epimorphism surjectivity property. By 12.20 we conclude that $\mathsf{M}$ is a discriminator variety. ∎

### v2

@@ 12. Structure theory
By symmetry it suffices to show that the statement holds for the case where $\phi=\mathsf{Ker}(g)$. Since $g\colon\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}\to\bm{C}$ is a homomorphism with $\bm{C}\in\mathsf{K}$, we have $\mathsf{Ker}(g)\in\mathsf{Con}_{\mathsf{K}}(\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*})$. Moreover, recall that $\bm{C}\in\textsf{K}_{\textsc{rfsi}}$ and that $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras by assumption. Therefore, from $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}/\mathsf{Ker}(g)\in\mathbb{I}\mathbb{S}(\bm{C})$ it follows that $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}/\mathsf{Ker}(g)$ is either trivial or belongs to $\mathsf{K}_{\textsc{rfsi}}$. Lastly, recall that $\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)^{*}\leqslant\bm{A}\times\bm{A}$ is a subdirect product with $\bm{A}\in\mathsf{K}$. As $\mathsf{K}$ is a relatively congruence distributive quasivariety by assumption, we can apply Corollary 2.18, obtaining the desired conclusion. ∎
By symmetry and Claim 12.16 we may assume that there exist $\eta_{1},\eta_{2}\in\mathsf{Con}_{\mathsf{K}}(\bm{A})$ such that $\mathsf{Ker}(g)=(\eta_{1}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$ and $\mathsf{Ker}(h)\in\{(\eta_{2}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)},(A^{2}\times\eta_{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}\}$. We have two cases depending on whether $\mathsf{Ker}(h)$ is $(\eta_{2}\times A^{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$ or $(A^{2}\times\eta_{2}){\upharpoonright}_{\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)}$.
@@
Now, recall that $\langle a,b\rangle\in\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)$. By Claim 12.17 we obtain $\langle a,b\rangle\in\eta_{1}$. Since $b\in A$, this implies $\langle\langle a,b\rangle,\langle b,b\rangle\rangle\in(\eta_{1}\times A^{2})$. On the other hand, from $a,b\in A$ and $\eta_{2}\in\mathsf{Con}(\bm{A})$ it follows that $\langle\langle a,b\rangle,\langle b,b\rangle\rangle\in(A^{2}\times\eta_{2})$. As $\langle a,b\rangle,\langle b,b\rangle\in\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(B)$, we obtain
@@
We begin by showing that $\mathsf{K}$ is a variety. As $\mathsf{K}$ is a quasivariety, by Proposition 2.5 it suffices to show that $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A})$ for every $\bm{A}\in\mathsf{K}$. To this end, consider $\bm{A}\in\mathsf{K}$ and $\theta\in\mathsf{Con}(\bm{A})$. We will show that $\theta=\mathsf{Cg}_{\mathsf{K}}^{\bm{A}}(\theta)$ which, in turns, implies $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{A})$, as desired. In view of Proposition 2.8, the congruence $\theta$ is the universe of a subalgebra $\bm{B}\leqslant\bm{A}\times\bm{A}$. Furthermore, for every $\langle a,b\rangle\in B=\theta$ we have $\langle a,a\rangle,\langle b,b\rangle\in\theta=B$ because $\theta$ is a congruence of $\bm{A}$. Therefore, we can apply Proposition 12.15, obtaining
@@
Let $\mathsf{M}$ be a congruence preserving Beth companion of $\mathsf{K}$. Since $\mathsf{M}$ is a pp expansion of $\mathsf{K}$ and $\mathsf{K}$ is a quasivariety by assumption, we obtain that $\mathsf{M}$ is also a quasivariety (see Theorem 10.3(ii)). In addition, as $\mathsf{M}$ is a Beth companion of $\mathsf{K}$ and $\mathsf{K}$ is a quasivariety, $\mathsf{M}$ has the strong epimorphism surjectivity property by Theorem 11.6. From the assumption that $\mathsf{K}$ is relatively congruence distributive and Theorem 12.8 it follows that $\mathsf{M}$ is relatively congruence distributive as well. Lastly, we will prove that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras. Consider $\bm{A}\leqslant\bm{B}\in\mathsf{M}_{\textsc{rfsi}}$ with $\bm{A}$ nontrivial. By Theorem 12.9 we have $\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$. Furthermore, $\bm{A}\leqslant\bm{B}$ implies $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{A}$ is nontrivial, so is $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$ and the assumption that $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras guarantee that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rfsi}}$. Together with $\bm{A}\in\mathsf{M}$, this allows us to apply Theorem 12.9, obtaining $\bm{A}\in\mathsf{M}_{\textsc{rfsi}}$. Hence, we conclude that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras.
Therefore, $\mathsf{M}$ is a quasivariety with the strong epimorphism surjectivity property that, moreover, is relatively congruence distributive and such that $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras. Thus, from Theorem 12.19 it follows that $\mathsf{M}$ is an arithmetical variety with the congruence extension property. Since $\mathsf{M}$ is a variety, we obtain $\mathsf{M}_{\textsc{fsi}}=\mathsf{M}_{\textsc{rfsi}}$. As $\mathsf{M}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras, we conclude that so is $\mathsf{M}_{\textsc{fsi}}$. ∎
@@
Let $\mathsf{K}$ be a relatively filtral quasivariety with the strong epimorphism surjectivity property. As $\mathsf{K}$ is relatively filtral, it is relatively congruence distributive and $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras (see, e.g., [24, Cor. 6.5(i, iv)]). Therefore, from Theorem 12.19 it follows that $\mathsf{K}$ is an arithmetical variety. Since $\mathsf{K}$ is a relatively filtral quasivariety, this yields that $\mathsf{K}$ is a congruence permutable filtral variety. As congruence permutable filtral varieties coincide with discriminator varieties (see, e.g., [18, 52]), we conclude that $\mathsf{K}$ is a discriminator variety. ∎
Furthermore, we recall that a join semilattice $\langle A;\lor\rangle$ is said to be dually Brouwerian when for all $a,b\in A$ there exists the smallest $c\in A$ such that $a\leqslant b\lor c$. Moreover, an element $a$ of a complete lattice $\bm{A}$ is compact when for every $X\subseteq A$ such that $a\leqslant\bigvee X$ there exists a finite $Y\subseteq X$ such that $a\leqslant\bigvee Y$. With every quasivariety $\mathsf{K}$ and algebra $\bm{A}$ we associate a join semilattice $\mathsf{Comp}_{\mathsf{K}}(\bm{A})$ whose universe is the set of compact elements of $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ and whose join operation ++ is defined by the rule $\theta+\phi=\mathsf{Cg}^{\bm{A}}_{\mathsf{K}}(\theta\cup\phi)$. Lastly, a member $\bm{A}$ of a quasivariety $\mathsf{K}$ is said to be simple relative to $\mathsf{K}$ when $\mathsf{Con}_{\mathsf{K}}(\bm{A})$ has exactly two elements, and we say that $\mathsf{K}$ is relatively semisimple when every member of $\mathsf{K}_{\textsc{rsi}}$ is simple relative to $\mathsf{K}$. When $\mathsf{K}$ is a variety, we simply say that $\bm{A}$ is simple and $\mathsf{K}$ semisimple. We will rely on the fact that a quasivariety $\mathsf{K}$ is relatively filtral if and only if it is relatively semisimple and $\mathsf{Comp}_{\mathsf{K}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{K}$ (see [24, Thm. 6.3] and [78, Thms. 5 & 8]). [^9] This description of filtrality originated in the context of varieties (see [51, 52]) and was later extended to quasivarieties in [24]. Bearing this in mind, we will now prove Corollary 12.11.
@@
Let $\mathsf{M}$ be a Beth companion of a relatively filtral quasivariety $\mathsf{K}$. Since every relatively filtral quasivariety has the relative congruence extension property (see, e.g., [24, Cor. 6.5(i)]), so does $\mathsf{K}$. As $\mathsf{M}$ is a pp expansion of $\mathsf{K}$, we conclude that $\mathsf{M}$ is congruence preserving by Theorem 12.4(ii).
We will show that $\mathsf{M}$ is also a relatively filtral quasivariety. The fact that $\mathsf{M}$ is a quasivariety is a consequence of Theorem 10.3(ii) and the assumption that $\mathsf{M}$ is a pp expansion of the quasivariety $\mathsf{K}$. Therefore, it only remains to prove that $\mathsf{M}$ is relatively filtral, i.e., that it is relatively semisimple and $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{M}$. To show that $\mathsf{M}$ is relatively semisimple, consider $\bm{A}\in\mathsf{M}_{\textsc{rsi}}$. Since $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$, we can apply Theorem 12.9, obtaining $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}_{\textsc{rsi}}$. As the quasivariety $\mathsf{K}$ is relatively semisimple (because it is relatively filtral by assumption), we obtain that $\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ has exactly two elements. Moreover, $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ because $\bm{A}\in\mathsf{M}$ and $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$. Therefore, $\mathsf{Con}_{\mathsf{M}}(\bm{A})$ has exactly two elements, whence $\bm{A}$ is simple relative to $\mathsf{M}$. Hence, we conclude that $\mathsf{M}$ is relatively semisimple, as desired. Next we prove that $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$ is dually Brouwerian for every $\bm{A}\in\mathsf{M}$. Consider $\bm{A}\in\mathsf{M}$ and observe that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ by Proposition 10.2. Since $\mathsf{M}$ is congruence preserving, we have $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$, whence $\mathsf{Comp}_{\mathsf{M}}(\bm{A})=\mathsf{Comp}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$. As $\mathsf{Comp}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ is dually Brouwerian (because $\mathsf{K}$ is relatively filtral and $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$), we conclude that so is $\mathsf{Comp}_{\mathsf{M}}(\bm{A})$. Thus, we conclude that $\mathsf{M}$ is a relatively filtral quasivariety.
Lastly, observe that $\mathsf{M}$ has the strong epimorphism surjectivity property because it is a Beth companion of the quasivariety $\mathsf{K}$ (see Theorem 11.6). Thus, $\mathsf{M}$ is a relatively filtral quasivariety with the strong epimorphism surjectivity property. By Corollary 12.20 we conclude that $\mathsf{M}$ is a discriminator variety. ∎